OpenStudy (anonymous):

What is the 9th term of the geometric sequence where a1 = –5 and a6 = –5,120?

OpenStudy (dumbcow):

\[a_{n} = a_{1}*r^{n-1}\] \[a_{6} = -5120 = -5*r^{6-1}\] \[r^{5} = \frac{5120}{5} = 1024\] \[r = 4\] \[\rightarrow a_{9} = -5*(4)^{9-1}\]

OpenStudy (anonymous):

no caculator do u know what it is?

OpenStudy (dumbcow):

there are calculators everywhere on your computer or online graphcalc.com